DeMey Corporation's selling price was $20 per unit. Fixed expenses totaled $54,000, variable expenses were $14 per unit, and the company reported a profit of $9,000 for the year. The break-even point for DeMey Corporation is:

8,500 units

10,500 units

9,000 units

4,500 units
Answer #1

Contribution margin=Sales-Variable cost  

=(20-14)=$6 per unit

Breakeven=Fixed expenses/Contribution margin

=54000/6

which is equal to

=9000 units